OpenStudy (gabylovesyou):

Which of the following is the slope between the two points (-5, -1) and (6, -4)? 3/11 -3/11 11/3 -11/3

OpenStudy (anonymous):

slope is the change in y over the change in x. work it out.

OpenStudy (anonymous):

11/3

OpenStudy (anonymous):

i think it's 9000

OpenStudy (anonymous):

let (x1,y1) = (-5,-1) (x2,y2) = (6,-4) use : \[m = \frac{ y2 - y1 }{ x2 - x1 }\] to find the slope m

OpenStudy (anonymous):

give the student the answer and she will solve the question, give the student the method, and she will solve all the questions

OpenStudy (gabylovesyou):

its C ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yea cuz (-4) - (-1) = 11

OpenStudy (anonymous):

use what i wrote to calculate it

OpenStudy (anonymous):

it's not C!!

OpenStudy (anonymous):

lol pete, let her work it out

OpenStudy (anonymous):

slope = rise/run = change in y/ change in x

OpenStudy (gabylovesyou):

its B

OpenStudy (gabylovesyou):

right?

OpenStudy (anonymous):

i think your totally right

OpenStudy (gabylovesyou):

yaay me :p lol

OpenStudy (anonymous):

yes.. B:) sorry for not being here